# Predictions of Stellar Occultations by Haumea and the Event of 4 May 2026

### 16 Mar 2026 v1

@@ Abstract
Context. Haumea is the third-largest of the five officially recognized dwarf planets and one of the four that reside in the trans-Neptunian region. It is among the most exotic bodies in the Solar System, with an exceptionally rapid rotation, a highly elongated triaxial shape, and a ring that orbits about three times more slowly than Haumea itself. Because of its large heliocentric distance, direct exploration by dedicated space missions is not feasible in the short term, so progress must rely on ground- and near-Earth facilities. Stellar occultations are among the most powerful tools to investigate trans-Neptunian objects.
Aims. We present new predictions of stellar occultations by Haumea and its ring for stars down to Gaia G = 21, and assess their scientific potential, with special emphasis on the 4 May 2026 event.
Methods. We computed occultation opportunities for the coming years and evaluated the 4 May 2026 geometry in detail, including Haumea’s rotation phase, known 3D shape, pole orientation, and sky-plane motion, to estimate the expected shadow-path width. Because the target star has a very large Gaia RUWE, we also carried out a dedicated reliability analysis, including speckle observations.
Results. We identify eleven valuable events through 2030. For 4 May 2026, we derive an expected sky-plane shadow width of 2224 $\pm$ 30 km, substantially larger than conservative nominal assumptions and therefore highly favorable for observations. Speckle imaging reveals a companion at $\sim$0.12 arcsec and $\Delta m\approx 3.1$; this companion is also expected to be occulted and shifts the nominal main-star path prediction on Earth by about 8 mas.
Conclusions. These results confirm the strong scientific return expected from coordinated observations of upcoming Haumea occultations, especially the 4 May 2026 event, and provide an updated framework to improve constraints on Haumea’s shape, density, ring properties, and environment.
Key Words.:
@@ 1 Introduction
But the 2017 occultation by Haumea also revealed that its size was significantly larger than previously thought. As a consequence, Haumea’s inferred density decreased from the then‑accepted value of around 2700 kg m-3, to approximately 2000 kg m-3, consistent with the densities of other large bodies such as Pluto and Triton (Ortiz et al., 2017). However, because this occultation provided only a single geometric cross‑section, assumptions had to be made regarding the ratio of Haumea’s two largest axes, using constraints derived from the amplitude of its rotational light curve. The precise relationship between light‑curve amplitude and axis ratio depends on the photometric model adopted to describe the surface scattering law. Although the pole orientation of the ring, assumed to be equatorial, could in principle allow reconstruction of the full 3‑D shape without additional assumptions, this requires accurate knowledge of Haumea’s rotational phase at the time of the occultation. Furthermore, the ring plane may precess slightly due to Haumea’s significant J2 component of its gravitational potential (because of Haumea’s elongated shape) or may exhibit small eccentricities, both of which complicate the determination of the ring pole. As a result, Haumea’s triaxial shape can still be refined, and the same is true for its density, through additional occultation observations.
An accurate density estimate is essential for modeling Haumea’s interior and determining whether it is in hydrostatic equilibrium, an issue that remains unresolved although a differentiated body may be a solution (Dunham et al., 2019). Besides, if a sufficient number of occultation chords are obtained in future events, Haumea’s topography could be reconstructed in detail, similarly to what has recently been achieved for the large TNO (307261) Máni, provisionally designated 2002 MS4 (Rommel et al., 2023). Moreover, much remains to be learned about the ring. For instance, if the ring particles are small, on the order of a few microns, the opacity can depend strongly on wavelength (Kalup and Kiss, 2025), allowing the particle‑size distribution to be constrained through multi‑wavelength occultation observations (e.g., Santos-Sanz et al., 2025). Additional faint satellites or narrower rings may be detected, and tighter upper limits on any possible atmosphere could be derived. The large satellite Hi‘iaka may also induce precession effects on the ring, which merit investigation.
@@ 2 Predictions
Predictions of stellar occultations by Haumea involving stars down to magnitude 21 in the Gaia G band were carried out for the interval 2024–2031 (inclusive). To do this, we extracted the positions of all Gaia DR3 stars located within the sky region traversed by Haumea through 2031. We then searched for close angular approaches between Haumea and those stars as seen from Earth, retaining only those cases with the potential to produce a ground‑track (shadow path) on any region of the Earth. The source of Haumea’s ephemerides was the JPL Horizons system. Several events with some observability potential were identified (see table 1). Maps showing the predicted shadow paths and auxiliary information for these events are provided in Appendix A (Figs. 414).
@@ 3 The 2026 May 4 Event
This detected companion affects the occultation prediction because the system’s photocenter is shifted by approximately 8 milliarcseconds toward the companion relative to the primary star’s nominal position in Gaia DR3. This is not a severe shift given the much larger angular size of the shadow. The northward star position is preferred, although the southward solution cannot be discarded because of the residual 180° ambiguity in the speckle reconstruction.
@@
To investigate the photometric variability we made observations with the Las Cumbres Observatory (LCO) 0.35m Delta Rho 350 telescopes equipped with QHY600 sCMOS cameras (see instrument details https://lco.global/observatory/instruments/qhy600-delta-rho-350/). We used the r_sloan filter for 5 nights to monitor the star brightness to look for potential short-term variability. The observations were carried out from 2025/05/17 to 2025/05/17 using 40s exposures. The Tenerife, Haleakala, MacDonald and Siding Spring sites could make observations. There was no continuous monitoring from the first night to the last night without gaps, but there was almost complete 24-h coverage in the first and last nights. Synthetic aperture photometry was carried out and instrumental magnitudes converted to the Johnson R system were obtained. Fig. 2 illustrates the time coverage and the derived magnitude. The errors of the individual measurements are only 0.01 mag on average. A Lomb-Scargle periodogram analysis shows a prominent peak at 6.52h with a confidence level well above the 99% level and the light curve folded to that period has a peak to valley amplitude of 0.013 mag (bottom panel of Fig. 2). This seems consistent with the possibility that the high RUWE of the star could come from variability in the secondary or in the primary.
@@
We also investigated whether the proper motion values in Gaia DR3 might be affected by the high RUWE and result in a poor propagation of the position of the star for the occultation time. However, the Gaia DR3 values are consistent with those listed in the HSOY catalog (Altmann et al., 2017), which is based mostly on PPMXL, combined with Gaia DR2 positions. HSOY lists proper motions of -7.724 mas/yr (RA) and -8.740 mas/yr (Dec), for the star, while GaiaDR3 gives -7.394 ± 0.170 mas/yr and -8.833 ± 0.139 mas/yr. Propagating the positions in the catalogs to the epoch of the occultation yields a difference of only 8 mas in declination (+14 40 25.612 for HSOY vs. +14 40 25.604 for Gaia DR3), while the right ascension values agree to the milliarcsecond level (14:40:58.4947 in both cases). Thus, even if Gaia’s proper motions were slightly biased, the impact on the prediction would be minimal.
@@
Another aspect to consider for this occultation is the rotation phase of Haumea. Because Haumea is highly elongated and rotates rapidly, its projected area changes substantially over a rotation cycle, and consequently the sky-plane shadow width also varies. Therefore, it is important to know Haumea’s rotational phase at the time of the occultation. We computed this using a light curve obtained in 2022 with the 1.5‑m telescope at Sierra Nevada Observatory, and then propagated the phase using the rotation period of 3.915341 ± 0.000005 hours (Ortiz et al., 2017). At the time of the occultation, the rotational phase will be only 25 degrees from maximum brightness (corresponding to maximum projected size). Using the spin‑axis orientation derived in Ortiz et al. (2017) and the 3D shape there, we estimate a shadow width of 2224 ± 30 km. This is a favorable result, as it indicates a much wider shadow path than the conservative nominal value of 1380 km used for initial predictions and highly enhances the detectability of the main‑body occultation (see Fig. 3).
@@
Another factor to consider is that Haumea possesses a ring system, meaning that the star’s occultation by the ring may be observable across a considerably wider region of Earth than the main‑body occultation itself (see Fig. 3).
@@
Regarding the effect of Haumea’s motion around the barycenter of the Haumea–Hi‘iaka system, we should point out that our predictions use the most recent JPL Horizons orbit and this corresponds to the barycenter of the system: Since most astrometric measurements assimilated into Horizons correspond to unresolved images of the Haumea system, the positions refer to the system's photocenter, which orbits the barycenter. Over long data arcs this motion averages out, so the fitted heliocentric orbit effectively represents the system barycenter. However, given current mass estimates (Proudfoot et al., 2024), the barycenter would lie roughly 6 mas away from Haumea toward Hi‘iaka. The resulting offset depends on Hi‘iaka’s orbital phase. For the May 4 event, Hi‘iaka is predicted to be at ($\DeltaRA, Δ\DeltaDec$) = (-0.03”, +1.02”), which corresponds to corrections of only +0.1 mas in right ascension and -4.7 mas in declination when converting barycentric coordinates to Haumea’s position. These corrections produce small changes (about 166 km cross-track) in the predicted ground track compared to the width of the path.
Note that observations of the event can be reported via the occultation portal (Kilic et al., 2022) at the following entry: https://opop.obspm.fr/create_report/2518/ where zoomable maps are also available.
@@ 4 Summary
Although the 180° ambiguity in the speckle position angle is not fully resolved, the northward companion star is currently preferred. In either case, the occultation can still be recorded from numerous sites, and an event by the fainter companion is expected most likely over northern Europe, with a secondary possible solution over southern Africa. Given the large angular extent of the ring system, the ring occultation will be observable over a wide fraction of the Earth. After the occultation is observed we will also be able to learn more about the situations that result in large RUWE in the Gaia DR3 catalog and we will be able to make progress on the knowledge of stellar binarity through the combination of Gaia data and stellar occultation results. We also presented other occultation opportunities, including visibility maps, to assist with detailed observation planning.
[…]
Appendix A Occultation maps
Figure 4: Stellar occultation prediction map and geometry for Haumea event 1 of 11. Time markers (first/middle/last): 20h 20m 40s / 20h 16m 30s / 20h 12m 30s UTC (50 markers). Note that the correct path prediction for this event is shown in Fig. 3 of the main paper
Figure 5: Stellar occultation prediction map and geometry for Haumea event 2 of 11. Time markers (first/middle/last): 2h 11m 40s / 2h 18m 20s / 2h 24m 50s UTC (80 markers).
Figure 6: Stellar occultation prediction map and geometry for Haumea event 3 of 11. Time markers (first/middle/last): 13h 53m 0s / 13h 59m 40s / 14h 6m 20s UTC (81 markers).
Figure 7: Stellar occultation prediction map and geometry for Haumea event 4 of 11. Time markers (first/middle/last): 15h 34m 10s / 15h 38m 10s / 15h 42m 10s UTC (50 markers).
Figure 8: Stellar occultation prediction map and geometry for Haumea event 5 of 11. Time markers (first/middle/last): 12h 17m 10s / 12h 13m 30s / 12h 9m 50s UTC (45 markers).
Figure 9: Stellar occultation prediction map and geometry for Haumea event 6 of 11. Time markers (first/middle/last): 1h 8m 30s / 1h 13m 0s / 1h 17m 20s UTC (54 markers).
Figure 10: Stellar occultation prediction map and geometry for Haumea event 7 of 11. Time markers (first/middle/last): 22h 33m 20s / 22h 37m 40s / 22h 42m 0s UTC (53 markers).
Figure 11: Stellar occultation prediction map and geometry for Haumea event 8 of 11. Time markers (first/middle/last): 10h 23m 50s / 10h 27m 30s / 10h 31m 10s UTC (45 markers).
Figure 12: Stellar occultation prediction map and geometry for Haumea event 9 of 11. Time markers (first/middle/last): 14h 21m 20s / 14h 15m 50s / 14h 10m 20s UTC (67 markers).
Figure 13: Stellar occultation prediction map and geometry for Haumea event 10 of 11. Time markers (first/middle/last): 19h 10m 30s / 19h 7m 10s / 19h 3m 50s UTC (41 markers).
Figure 14: Stellar occultation prediction map and geometry for Haumea event 11 of 11. Time markers (first/middle/last): 14h 24m 30s / 14h 21m 0s / 14h 17m 40s UTC (42 markers).
Acknowledgements.

### 16 Apr 2026 v2

@@ Abstract
Haumea is the third-largest of the five officially recognized dwarf planets and one of the four that reside in the trans-Neptunian region. It is among the most exotic bodies in the Solar System. Because of its large distance, direct exploration by space missions is not feasible in the short term, so progress must rely on ground- and near-Earth facilities. Stellar occultations are among the most powerful tools to investigate trans-Neptunian objects. We present predictions of occultations by Haumea in the coming years for stars down to Gaia G = 21. We identify eleven relevant events through 2030. We analysed in detail the geometry of the 4 May 2026 event, including Haumea’s rotation phase, known 3D shape, pole orientation, and sky-plane motion. We derive a sky-plane shadow width of 2224 $\pm$ 30 km, substantially larger than nominal assumptions and therefore highly favorable for observations. Given the star’s large RUWE of 6.6, raising concerns about the detectability of the event, we also performed a reliability analysis. Speckle imaging reveals a companion at $\sim$0.12 arcsec and $\Delta m\approx 3.1$; this companion is also expected to be occulted, and shifts the nominal main-star path prediction on Earth by about 8 mas.
keywords:
@@ 1 Introduction
But the 2017 occultation by Haumea also revealed that its size was significantly larger than previously thought. As a consequence, Haumea’s inferred density decreased from the then‑accepted value of around 2700 kg m-3, to approximately 2000 kg m-3, consistent with the densities of other large bodies such as Pluto and Triton (Ortiz et al., 2017). However, because this occultation provided only a single geometric cross‑section, assumptions had to be made regarding the ratio of Haumea’s two largest axes, using constraints derived from the amplitude of its rotational light curve. The precise relationship between light‑curve amplitude and axis ratio depends on the photometric model adopted to describe the surface scattering law. Although the pole orientation of the ring, assumed to be equatorial, could in principle allow reconstruction of the full 3‑D shape without additional assumptions, this requires accurate knowledge of Haumea’s rotational phase at the time of the occultation (e.g., Dunham et al., 2019). Furthermore, the ring plane may precess slightly due to Haumea’s significant J2 component of its gravitational potential (because of Haumea’s elongated shape) or may exhibit small eccentricities, both of which complicate the determination of the ring pole. Recent dynamical modelling suggests that long-term precession of Haumea’s ring plane should be very small: when the effects of the body’s triaxial figure, satellite perturbations, and collisional damping are taken into account, the ring is expected to remain close to the equatorial plane, although non-negligible particle eccentricities may still be sustained (Marzari, 2020). As a result, Haumea’s triaxial shape can still be refined, and the same is true for its density, through additional occultation observations.
An accurate density estimate is essential for modeling Haumea’s interior and determining whether it is in hydrostatic equilibrium, an issue that remains unresolved although a differentiated body may be a solution (Dunham et al., 2019). Besides, if a sufficient number of occultation chords are obtained in future events, Haumea’s topography could be reconstructed in detail, similarly to what has recently been achieved for the large TNO (307261) Máni, provisionally designated 2002 MS4 (Rommel et al., 2023). Moreover, much remains to be learned about the ring. For instance, if the ring particles are small, on the order of a few microns, the opacity can depend strongly on wavelength (Kalup and Kiss, 2025), allowing the particle‑size distribution to be constrained through multi‑wavelength occultation observations (e.g., Santos-Sanz et al., 2025). Additional faint satellites or narrower rings may be detected, and tighter upper limits on any possible atmosphere could be derived. The large satellite Hi‘iaka may also induce precession effects on the ring, which merit investigation. Marzari (2020) showed that Hi‘iaka likely does not have a strong effect on the nodal precession of the rings since it is coplanar with the rings to within $\sim$1. Namaka, although smaller, can induce a larger nodal precession effect because of its higher $\sim$13 inclination. On the other hand, apsidal precession may be dominated by Hi‘iaka.
@@ 2 Predictions
Predictions of stellar occultations by Haumea involving stars down to magnitude 21 in the Gaia G band were carried out for the interval 2024–2031 (inclusive). To do this, we extracted the positions of all Gaia DR3 stars located within the sky region traversed by Haumea through 2031. We then searched for close angular approaches between Haumea and those stars as seen from Earth, retaining only those cases with the potential to produce a ground‑track (shadow path) on any region of the Earth. The source of Haumea’s ephemerides was the JPL Horizons system. Several events with some observability potential were identified (see table 1). Maps showing the predicted shadow paths and auxiliary information for these events are provided as Supplementary Material (occultation maps for events 1–11).
@@ 3 The 2026 May 4 Event
This detected companion affects the occultation prediction because the system’s photocenter is shifted by approximately 8 milliarcseconds toward the companion relative to the primary star’s nominal position in Gaia DR3. This is not a severe shift given the much larger angular size of the shadow. Bispectral analysis of the phase information contained in the images (Horch et al., 2011) can be used to select the correct quadrant; in our case it favoured the northern location for the companion star, although the southward solution cannot be completely discarded.
@@
To investigate the photometric variability we made observations with the Las Cumbres Observatory (LCO) 0.35m Delta Rho 350 telescopes equipped with QHY600 sCMOS cameras (see instrument details https://lco.global/observatory/instruments/qhy600-delta-rho-350/). We used the r_sloan filter for 5 nights to monitor the star brightness to look for potential short-term variability. The observations were carried out from 2025/05/17 to 2025/05/22 using 40s exposures. The Tenerife, Haleakala, MacDonald and Siding Spring sites could make observations. There was no continuous monitoring from the first night to the last night without gaps, but there was almost complete 24-h coverage in the first and last nights. Synthetic aperture photometry was carried out and instrumental magnitudes converted to the Johnson R system were obtained. Fig. 2 illustrates the time coverage and the derived magnitude. The errors of the individual measurements are only 0.01 mag on average. A Lomb-Scargle periodogram analysis shows a prominent peak at 6.52h with a confidence level well above the 99% level and the light curve folded to that period has a peak to valley amplitude of 0.013 mag (bottom panel of Fig. 2). This seems consistent with the possibility that the high RUWE of the star could come from variability in the secondary or in the primary.
@@
We also investigated whether the proper motion values in Gaia DR3 might be affected by the high RUWE and result in a poor propagation of the position of the star for the occultation time. However, the Gaia DR3 values are consistent with those listed in the "Hot Stuff for One Year" (HSOY) catalog (Altmann et al., 2017), which is based mostly on the PPMXL catalog (Roeser et al., 2010), combined with Gaia DR2 positions. HSOY lists proper motions of -7.724 mas/yr (RA) and -8.740 mas/yr (Dec), for the star, while GaiaDR3 gives -7.394 ± 0.170 mas/yr and -8.833 ± 0.139 mas/yr. Propagating the positions in the catalogs to the epoch of the occultation yields a difference of only 8 mas in declination (+14 40 25.612 for HSOY vs. +14 40 25.604 for Gaia DR3), while the right ascension values agree to the milliarcsecond level (14:40:58.4947 in both cases). Thus, even if Gaia’s proper motions were slightly biased, the impact on the prediction would be minimal.
@@
Another aspect to consider for this occultation is the rotation phase of Haumea. Because Haumea is highly elongated and rotates rapidly, its projected area changes substantially over a rotation cycle, and consequently the sky-plane shadow width also varies. Therefore, it is important to know Haumea’s rotational phase at the time of the occultation. We computed this using a light curve obtained in 2022 with the 1.5‑m telescope at Sierra Nevada Observatory, and then propagated the phase using the rotation period of 3.915341 ± 0.000005 hours (Ortiz et al., 2017). At the time of the occultation, the rotational phase will be only 25 degrees from maximum brightness (corresponding to maximum projected size). Using the spin‑axis orientation derived in Ortiz et al. (2017) and the 3D shape there, we estimate a shadow width of 2224 ± 30 km. This is a favorable result, as it indicates a much wider shadow path than the conservative nominal value of 1380 km used for initial predictions and highly enhances the detectability of the main‑body occultation (see Fig. 3). A recent study (Staelen et al., 2026) also suggests that Haumea may possess an equatorial deformation large enough to be detectable in forthcoming occultations, which further strengthens the scientific interest of well-sampled events such as that of 2026 May 4.
@@
Another factor to consider is that Haumea possesses a ring system, meaning that the star’s occultation by the ring may be observable across a considerably wider region of Earth than the main‑body occultation itself (see Fig. 3). The shadow path corresponding to the companion star in its preferred position for both the main body and the ring is shown in Fig. 4.
@@
Figure 4: Same as Fig. 3 but for the companion star in its preferred position. The arrow indicates the motion of the shadow.
Regarding the effect of Haumea’s motion around the barycenter of the Haumea–Hi‘iaka system, we should point out that our predictions use the most recent JPL Horizons orbit and this corresponds to the barycenter of the system: Since most astrometric measurements assimilated into Horizons correspond to unresolved images of the Haumea system, the positions refer to the system's photocenter, which orbits the barycenter. Over long data arcs this motion averages out, so the fitted heliocentric orbit effectively represents the system barycenter. However, given current mass estimates (Proudfoot et al., 2024), the barycenter would lie roughly 6 mas away from Haumea toward Hi‘iaka. The resulting offset depends on Hi‘iaka’s orbital phase. For the May 4 event,according to the Miriade service [^1] https://vo.imcce.fr/webservices/miriade, Hi‘iaka is predicted to be at ($\DeltaRA, Δ\DeltaDec$) = (-0.03", +1.02"), which corresponds to corrections of only +0.1 mas in right ascension and -4.7 mas in declination when converting barycentric coordinates to Haumea’s position. These corrections produce small changes (about 166 km cross-track) in the predicted ground track compared to the width of the path. Note that both Hi‘iaka and Namaka are angularly so far from Haumea at the date of the occultation that the corresponding shadow paths lie outside the Earth.
Observations of the event can be reported via the occultation portal (Kilic et al., 2022) at the following entry: https://opop.obspm.fr/create_report/2518/ where zoomable maps are also available.
@@ 4 Summary
Although the 180° ambiguity in the speckle position angle is not fully resolved, the northward companion star is currently preferred. In either case, the occultation can still be recorded from numerous sites, and an event by the fainter companion is expected most likely over southern Africa, with a secondary possible solution over northern Europe. Given the large angular extent of the ring system, the ring occultation will be observable over a wide fraction of the Earth. After the occultation is observed we will also be able to learn more about the situations that result in large RUWE in the Gaia DR3 catalog and we will be able to make progress on the knowledge of stellar binarity through the combination of Gaia data and stellar occultation results. We also presented other occultation opportunities, including visibility maps, to assist with detailed observation planning.
Supplementary material
Supplementary material is available online as a single PDF file (supplementary.pdf) containing the occultation prediction maps and geometry plots for events 1–11.
Acknowledgements
@@
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.